OpenStudy (anonymous):

Determine Which Ordered Pair Is A Solution of Y=2x-3. 1. (-2, 7) 2. (0, 3) 3. (-4, -11) 4. (5, -7)

OpenStudy (texaschic101):

ur gonna have to sub in ur answer choices into original eq and see which one is true

OpenStudy (texaschic101):

i will do first one... y = 2x - 3....(- 2,7) 7 = 2(- 2) - 3 7 = - 4 - 3 7 = - 7...not true..this is not the solution now u try the other answer choices

rebeccaxhawaii (rebeccaxhawaii):

plug your x's and y's in and if its a true statement then its a solution

OpenStudy (luckyness):

@rebeccaxhawaii Is correct. You need to plug in your x and y's.